OpenStudy (anonymous):

when you do this problem can you not only give me the answer but how you got it to please The equation of Circle P is ( x − 3)2 + (y + 2)2 = 34 and the equation of circle Q is (x − 4)2 + (y + 5)2 = 52. What is the distance between the centers of these circles?

OpenStudy (anonymous):

deja vu all over again. the center of the first one is \[(3,-2)\] with your eyeballs. likewise the center of the second one is \[(4,-5)\]

OpenStudy (anonymous):

now compute the distance between them via \[\sqrt{(3-4)^2+(-2+5)^2}\] \[\sqrt{1+9}\] \[\sqrt{10}\]

OpenStudy (anonymous):

thank you